Need help completing the ordered pairs (4, ?) and (?, 5)
that satisfy the equation 3x – 6y = 7.
here is my try: 3x-6y=7
/3 /3
6y/3
y=2?
Need help solving this and finding the ordered pairs. Thanks

Answer by rfer(16322)

3x-6y=7
3(4)-6y=7
12-6y=7
-6y=-5
y=5/6
(4,5/6)
------------
3x-6(5)=7
3x-30=7
3x=37
x=12 1/3
(12 1/3,5)